# LLMs as Planning Modelers: A Survey for Leveraging Large Language Models to Construct Automated Planning Models

### v1

@@ 3.3 Model Benchmarks
With the proliferation of related techniques to convert NL to PDDL, we are seeing an ever-increasing set of related methods. To bring them together under a single computational umbrella, and beyond just relating the work together conceptually as we have done thus far in this survey, we created a unified platform: L2P  [^1] Code made publicly at: https://github.com/AI-Planning/l2p, which re-implements landmark papers covered in this survey. This Python library is open source and has the capability of encapsulating the generalized version of the proposed “LLM-Modulo” framework (Kambhampati et al. 2024), which emphasizes soundness guarantees through iterative plan refining via external verifiers. While L2P embodies the core principles of the LLM-Modulo framework, which advocates for LLMs autonomously generating plan candidates themselves, it shifts the focus by facilitating the creation of PDDL files using LLMs – aligning with this paper’s paradigm. This approach allows for the integration of external verifiers and feedback critics, enabling users to modify and refine the extracted models. L2P offers three major benefits:
@@
(iii) Autonomous Capability: supports a fully autonomous pipeline, reducing the need for manual authoring.
Appendix B demonstrates several usage examples of L2P. To demonstrate the flexibility of the framework, we re-implemented several of the landmark papers covered in this survey, and Appendix C provides potential works that can be recreated by L2P at the time of writing. We hope that the community contributes to the L2P framework as a repository of existing advancements in LLM model acquisition and relevant papers, ensuring that users have access to the most current research and tools under a common framework for fair comparison.
4.1 Paper Reconstructions
L2P can recreate and encompass previous frameworks for converting natural language to PDDL, serving as a comprehensive foundation that integrates past approaches. The L2P GitHub contains multiple paper reconstructions as examples of how existing methods can be implemented and compared within a unified system, demonstrating L2P’s flexibility and effectiveness in standardizing diverse NL-to-PDDL techniques. An example of Guan et al. (2023) ”action-by-action” algorithm can be found in Figure 3; predicate output can be found in Appendix B Figure 7.
1import os
2from l2p import *
3
4def run_aba_alg(model: LLM, action_model,
5 domain_desc, hierarchy, prompt, max_iter: int=2
@@ 4.1 Paper Reconstructions
7actions = list(action_model.keys())
8pred_list = []
9
10for _ in range(max_iter):
@@
14 if len(pred_list) == 0:
15 prompt = prompt.replace(’{predicates}’,
[…]
17 else: res = ""
18 for i, p in enumerate(pred_list):
[…]
20 prompt = prompt.replace(’{predicates}’, res)
21 # extract pddl action and predicates (L2P)
22 pddl_action, new_preds, response = (
23 builder.extract_pddl_action(
@@
29 action_list=action_list,
@@
31 types=hierarchy["hierarchy"]
@@
38return pred_list, action_list
39
40if __name__ == "__main__":
41 builder = DomainBuilder() # domain build class
42 # load in assumption files
43 base_path=’paper_reconstructions/llm+dm/prompts/’
44 action_model=load_file(
45 f’{base_path}action_model.json’)
46 domain_desc=load_file(
47 f’{base_path}domain_desc.txt’)
48 hier=load_file(
49 f’{base_path}hierarchy_requirements.json’)
50 prompt=load_file(f’{base_path}pddl_prompt.txt’)
51 # initialise LLM engine (OpenAI in this case)
@@
54
55 # run "action-by-action" algorithm
56 pred, action = run_aba_alg(model=llm, action_model,
57 domain_desc, hier, prompt)
Figure 3: Shortened L2P reconstruction of “action-by-action algorithm” (Guan et al. 2023). This iterates through each action description and generates its PDDL specifications while maintaining a dynamically generating predicate list.
5 Discussion
Revisiting RQ1: While frameworks can generate parsable and solvable PDDL files, it remains uncertain if these specifications align with human goals. Simple domains like Blocksworld are easier to verify. Still, scaling complex domains require users to understand how LLMs generate these specifications, emphasizing the need for explainable planning for robust, transparent, and correctable outputs (Zuo et al. 2024). Corrective feedback loops have demonstrated significant improvements in resolving failing actions preconditions (Raman et al. 2024), or re-planning in case of unexpected failures during the execution of the plan (Joublin et al. 2023; Raman et al. 2022). Ensuring alignment with user goals involves breaking down PDDL model construction into pre-processing steps with human-in-the-loop feedback, as seen in frameworks like (Kelly et al. 2023). Very reminiscent of the “critics” process in the LLM-Modulo framework (Kambhampati et al. 2024), setting up a sort of external verifier checklist and using LLMs to provide feedback, is demonstrated by Gestrin, Kuhlmann, and Seipp (2024) and Smirnov et al. (2024). A strong idea is analyzing the semantic correctness of plans generated and using that as feedback to refine the LLM-generated PDDL specifications (Sakib and Sun 2024). Additionally, intermediate representation (i.e. ASP, Python, JSON) methods that are easier for LLMs to process before converting to PDDL (Agarwal and Sreepathy 2024; Smirnov et al. 2024) can also enhance accuracy. However, over time, LLMs are likely to improve, with fewer syntax errors and more logical semantic outputs as they train on more PDDL knowledge.
Revisiting RQ2: LLMs have demonstrated that they are significantly sensitive to prompting raising questions about whether they are better off functioning as machine translators or generators. Liu et al. (2023a) demonstrate that highly explicit descriptions improve translation accuracy, while Gestrin, Kuhlmann, and Seipp (2024) and Smirnov et al. (2024) leverage minimal descriptions, relying on LLMs’ internal world knowledge to enrich outputs. Huang, Lipovetzky, and Cohn (2024), Liu et al. (2023a), and Guan et al. (2023) recognize that specifying a precise predicate set in natural language is crucial and addresses the common problem of evaluating across different methods. The challenge of operating with minimal to no textual guidance beyond the initial task prompt becomes significantly more complex in online planning contexts. This underscores the importance of standardizing prompt granularity for initial generation and iterative feedback (Liu et al. 2024b). Nabizada et al. (2024) provides a promising organized and standardized paradigm of automated generated PDDL descriptions that can be applied to LLMs.
5.1 Literature Scope and Additional Outlooks
Our survey examines nearly 80 works, with over 50 integrating LLMs into the workflow for generating AP specifications. These frameworks span task (30 papers), domain (15 papers), and hybrid (15 papers) model generation. Strong emphasis on task modeling stems from the fact that many systems, already operate within well-defined domains–such as robotics. However, advancing toward more sophisticated and adaptable systems–such as dialogue agents–requires a more holistic approach, where both task objectives and domain constraints are dynamically constructed. We encourage researchers to use this survey as a foundation for exploring beyond task modeling and toward integrated domain and hybrid specification approaches, enabling more robust and scalable end-to-end planning.
6 Conclusion
The issue of extracting these planning models has long been recognized as a major barrier to the widespread adoption of planning technologies, a topic that has been explored in depth within the knowledge acquisition literature in AP for decades (Vallati and Kitchin 2020; Hendler, Tate, and Drummond 1990). Even with the emergence of LLMs, this remains a persistent challenge, introducing a new suite of obstacles. In this survey, we examine nearly 80 scholarly articles that propose their frameworks and some other subsidiary works delegating model acquisition tasks to LLMs. By identifying the research distribution and gaps within these categories, we aim to provide a higher generalization from each framework’s methodologies into broader aspects for future architectures. Additionally, we hope researchers can apply these methodologies to more advanced planning languages with the support of our L2P library–that can be found here: https://github.com/AI-Planning/l2p.
Limitations
This survey has two primary limitations. First, regarding its scope, our focus is limited to PDDL construction frameworks and related papers. Techniques remain largely unexplored in this context, and LLM capabilities in planning are still in their early stages. Due to page space constraints, we provide only a brief overview of each work rather than an exhaustive technical analysis. Additionally, our study primarily draws works published in ACL, ACM, AAAI, NeurIPS, ICAPS, COLING, CoRR, ICML, ICRA, EMNLP, and arXiv, so there is a possibility that we may have missed relevant research from other venues. Secondly, our L2P library currently supports only basic PDDL extraction tools for fully observable deterministic planning, and does not yet include tools for areas such as conditional, temporal, or numeric planning. We plan to expand the library to cover a broader range of PDDL applications, aiming to further research into the challenges LLMs encounter in these areas.
Ethics Statement
Appendix A Additional Information on PDDL
A.1 Why PDDL?
PDDL, as a declarative language, differs fundamentally from imperative languages like Python in how problems are expressed and solved. Declarative programming specifies what needs to be achieved rather than how to achieve it, leaving execution to an external planner. In contrast, imperative programming defines explicit step-by-step instructions, requiring precise control over execution flow. While LLMs struggle with logical reasoning in imperative programming due to its sequential dependencies, they can perform well with declarative representations like PDDL. Using LLMs for PDDL model construction is not traditional code generation but rather knowledge structuring—organizing states, actions, and constraints into a formalized model.
A.2 PDDL Example
Demonstrated below is the Blocksworld PDDL domain file $\mathbb{DF}$:
1(define (domain blocksworld)
2 (:requirements :strips)
3 (:predicates (clear ?x) (on-table ?x) (arm-empty) (holding ?x) (on ?x ?y))
4 (:action pickup
5 :parameters (?ob)
6 :precondition (and (clear ?ob) (on-table ?ob) (arm-empty))
7 :effect (and (holding ?ob) (not (clear ?ob)) (not (on-table ?ob)) (not (arm-empty)))
8 ))
1(define (problem blocksworld-problem)
2 (:domain blocksworld)
3 (:objects A B C) ; Blocks
4 (:init (ontable A) (ontable B) (on C A) (clear B) (clear C)) ; Initial state
5 (:goal (and (on A B) (on B C)))) ; Goal state
1Unstack C from A
2Put C on table
3Pick up A
4Stack A on B
5Pick up B
6Stack B on C
[…]
5
Appendix B (L2P) Framework
B.1 General Library Overview
L2P provides a comprehensive suite of tools for PDDL model creation and validation. The Builder classes enable users to prompt the LLM to generate essential components of a PDDL domain, such as types, predicates, and actions, along with their parameters, preconditions, and effects. It also supports task specification, including objects, initial, and goal states that correspond to the given domain. L2P features a customized Feedback Builder class that incorporates both LLM-generated feedback and human input, or a combination of the two. Additionally, the library includes a syntax validation tool that detects common PDDL syntax errors, using this feedback to improve the accuracy of the generated models—as illustrated in Figure 6, which shows how LLM feedback can be used to refine a PDDL problem specification. L2P is now publicly available for pip installation and can be found at: https://github.com/AI-Planning/l2p
B.2 L2P Usage Example
1 import os
2 from l2p import *
3
4 domain_builder = DomainBuilder() # initialize Domain Builder class
5
6 # REPLACE WITH OWN API KEY
7 api_key = os.environ.get(’OPENAI_API_KEY’)
8 llm = OPENAI(model="gpt-4o-mini", api_key=api_key)
9
@@ B.2 L2P Usage Example
13 extract_predicates_prompt = load_file(f’{base_path}extract_predicates.txt’)
@@
18 predicates, llm_output = domain_builder.extract_predicates(
@@
21 prompt_template=extract_predicates_prompt,
22 types=types,
23 nl_actions={action[’action_name’]: action[’action_desc’]}
@@
34 (holding ?a - arm ?b - block) ; true if the arm ?a is holding the block ?b
35 (on_top ?b1 - block ?b2 - block) ; true if the block ?b1 is on top of the block ?b2
36 (clear ?b - block) ; true if the block ?b is clear (no block on top of it)
37 (on_table ?b - block) ; true if the block ?b is on the table
38 (empty ?a - arm) ; true if the arm ?a is empty (not holding any block)
@@
2 from l2p import *
3
@@
10 extract_task_prompt = load_file(r’tests/usage/prompts/problem/extract_task.txt’)
@@
16 objects, initial_states, goal_states, llm_response = task_builder.extract_task(
@@
19 prompt_template=extract_task_prompt,
@@
23
24 # format key info into PDDL strings
25 objects_str = task_builder.format_objects(objects)
26 initial_str = task_builder.format_initial(initial_states)
27 goal_str = task_builder.format_goal(goal_states)
28
@@
31 domain="blocksworld",
32 problem="blocksworld_problem",
33 objects=objects_str,
34 initial=initial_str,
35 goal=goal_str)
36
@@
44 blue_block - object
45 red_block - object
46 yellow_block - object
47 green_block - object
@@
54 (clear blue_block) (clear yellow_block) (clear green_block)
@@
57 (and (on_top red_block green_block))
@@
2 from l2p import *
@@
15 objects, initial, goal, feedback_response = feedback_builder.task_feedback(
@@
21 types=types,
22 llm_response=llm_response)
@@
26 ### LLM OUTPUT
27 FEEDBACK:

### v2

@@ 4 Language-to-Plan (L2P)
With the proliferation of related techniques to convert NL to PDDL, we are seeing an ever-increasing set of related methods. To bring them together under a single computational umbrella, and beyond just relating the work together conceptually as we have done thus far in this survey, we created a unified platform: L2P  [^1] Code made publicly at: https://github.com/AI-Planning/l2p, which re-implements landmark papers covered in this survey. This Python library is open source and has the capability of encapsulating the generalized version of the proposed “LLM-Modulo” framework (Kambhampati et al., 2024), which ensures soundness via iterative plan refinement with external verifiers, shifting focus from direct planning to PDDL generation with integrated verifiers and user-guided refinement through complete, planner-executable specifications. L2P offers three major benefits:
@@
(iii) Autonomous Capability: supports fully autonomous end-to-end pipelines, reducing the need for manual authoring.
Appendix C demonstrates examples of L2P usage. To demonstrate the flexibility of the framework, L2P re-implements some key papers covered in this survey (refer to Appendix A). We hope to maintain the L2P framework as a repository of existing advancements in LLM model acquisition and relevant papers, ensuring that users have access to the most current research and tools.
1def run_aba_alg(
2 model, action_model, domain_desc, hier, prompt, max_iter: int=2
@@
4
5 actions = list(action_model.keys())
6 pred_list = []
7 for _ in range(max_iter):
@@
10 # extract action/predicates (L2P)
11 pddl_action, new_preds, _, _ = (
12 builder.formalize_pddl_action(
@@
18 types=hier["hierarchy"],
@@
20 extract_new_preds=True
@@
26 return pred_list, action_list
Figure 3: A shortened L2P reconstruction of the ’action-by-action algorithm’ (Guan et al., 2023), which iteratively generates PDDL actions while updating a dynamic predicate list. Output found in Figure 8.
5 Discussion
Revisiting RQ1: While frameworks can generate parsable and solvable PDDL files, it remains uncertain if these specifications align with human goals. Simple domains like Blocksworld are easier to verify. Still, scaling complex domains requires users to understand how LLMs generate these specifications, emphasizing the need for explainable planning to yield robust, transparent, and correctable outputs (Zuo et al., 2024). Corrective feedback loops notably improve failure handling, such as resolving action precondition errors (Raman et al., 2024), or re-planning in case of unexpected failures during plan execution (Joublin et al., 2023; Raman et al., 2022). Ensuring alignment with user goals involves breaking down PDDL model construction into pre-processing steps with human-in-the-loop feedback (Kelly et al., 2023). Very reminiscent of the “critics” process in the LLM-Modulo framework (Kambhampati et al., 2024), setting up a sort of external verifier checklist and using LLMs to provide feedback is demonstrated by Gestrin et al. (2024) and Smirnov et al. (2024). A potential idea is analyzing the semantic correctness of plans generated and using that as feedback to refine the LLM-generated PDDL specifications (Sakib and Sun, 2024). Additionally, intermediate representation (i.e. ASP, Python, JSON) that are easier for LLMs to process before converting to PDDL (Agarwal and Sreepathy, 2024; Smirnov et al., 2024) can also enhance accuracy.
Revisiting RQ2:
LLMs have demonstrated that they are significantly sensitive to prompting—raising questions about whether they are better off functioning as translators or generators. Liu et al. (2023a) demonstrate that highly explicit descriptions improve translation accuracy, while Gestrin et al. (2024) leverage minimal descriptions, relying on LLMs’ internal world knowledge to enrich outputs; however, this excess freedom often leads to inconsistent or inexecutable domain models. Huang et al. (2024b); Liu et al. (2023a); Guan et al. (2023) recognize that specifying a precise predicate set in NL is crucial and addresses the common problem of evaluating across different methods. The challenge of operating with minimal to no textual guidance beyond the initial task prompt underscores the importance of standardizing prompt granularity for initial generation and iterative feedback (Liu et al., 2024b). Nabizada et al. (2024) provide a promising, organized, and standardized paradigm for automatically generating PDDL descriptions that can be applied to LLMs.
6 Conclusion
Extracting planning models has long been recognized as a major barrier to the widespread adoption of planning technologies (Vallati and Kitchin, 2020; Hendler et al., 1990). Even with the emergence of LLMs, this remains a persistent challenge, introducing a new suite of obstacles. In this survey, we examine nearly 80 scholarly articles that propose their frameworks and some other subsidiary works delegating model acquisition tasks to LLMs. By identifying the research distribution and gaps within these categories, we aim to provide a higher generalization from each framework’s methodologies into broader aspects for future architectures. Additionally, we hope researchers can apply these methodologies to more advanced planning languages with the support of our L2P library.
Limitations
This survey has two primary limitations. First, regarding its scope, our focus is limited to PDDL construction frameworks and related papers. Techniques remain largely unexplored in this context, and LLM capabilities in planning are still in their early stages. Due to page space constraints, we provide only a brief overview of each work rather than an exhaustive technical analysis. Additionally, our study primarily draws works published in ACL, ACM, AAAI, NeurIPS, ICAPS, COLING, CoRR, ICML, ICRA, EMNLP, and arXiv, so there is a possibility that we may have missed relevant research from other venues. Secondly, our L2P library currently supports only basic PDDL extraction tools for fully observable deterministic planning, and does not yet include tools for areas such as temporal planning. We plan to expand the library to cover a broader range of PDDL applications, aiming to further research into the challenges LLMs encounter in these areas.
Ethics Statement
Appendix A Paper Overview
Below is a quick summary of core model generation frameworks in this survey. Papers are sorted by planning specification coverage, followed by publication year, and then alphabetically by author.
WORKS PDDL SPEC. GUIDANCE
Paper Framework Init. Goal Preds. Act. Prompt Feedback Human Int.
(Gestrin et al., 2024) NL2PLAN $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Heavy (Few-shot+CoT) Custom validator + LLM Optional
(Liu et al., 2024b) DELTA $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Medium (Few-shot) None None
(Mahdavi et al., 2024) LLM+EW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Light Env. (EW Metric) + LLM None
(Smirnov et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Light (JSON interm.) Custom validator + LLM Optional
(Sakib and Sun, 2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Medium (Few-shot) None None
(Ye et al., 2024) MORPHeus $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Heavy (Few-shot+CoT) NL Human feedback True
(Kelly et al., 2023) TABA $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Light (One-shot) Glaive validator + LLM Optional
(Ying et al., 2023) NIPE $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Medium (Few-shot) None None
(Zhou et al., 2023) ISR-LLM $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Medium (Few-shot) None None
(Han et al., 2024) InterPret $\times$ $\checkmark$ $\checkmark$ $\checkmark$ Light (CoT) NL Human feedback True
(Liu et al., 2024c) $\times$ $\checkmark$ $\checkmark$ $\checkmark$ Medium (unknown) Scene graph True
(Hu et al., 2025) TEXT2WORLD $\times$ $\times$ $\checkmark$ $\checkmark$ Heavy (Few-shot+Exp.) Tarski validator + LLM None
(Sinha, 2024) $\times$ $\times$ $\checkmark$ $\checkmark$ Heavy (Few-shot+Exp.) None None
(Sikes et al., 2024b) $\times$ $\times$ $\checkmark$ $\checkmark$ Medium (Few-shot) None None
(Guan et al., 2023) LLM+DM $\times$ $\times$ $\checkmark$ $\checkmark$ Heavy (Few-shot+CoT+Exp.) Custom validator + LLM None
(Chen et al., 2024) LASP $\times$ $\times$ $\times$ $\checkmark$ Medium (High Exp.) LLM None
(Huang et al., 2024b) $\times$ $\times$ $\times$ $\checkmark$ Medium (Few-shot) VSCode-PDDL validator None
(Liu et al., 2024a) BLADE $\times$ $\times$ $\times$ $\checkmark$ Heavy Human demonstration True
(Oates et al., 2024) CLLaMP $\times$ $\times$ $\times$ $\checkmark$ Medium (Few-shot) None None
(Oswald et al., 2024) NL2PDDL $\times$ $\times$ $\times$ $\checkmark$ Medium (Few-shot) None None
(Wong et al., 2023) ADA $\times$ $\times$ $\times$ $\checkmark$ Medium (Few-shot+Exp.) None None
(Zhang et al., 2024b) PROC2PDDL $\times$ $\times$ $\times$ $\checkmark$ Heavy (Few-shot) None None
(Athalye et al., 2024) pix2pred $\times$ $\times$ $\checkmark$ $\times$ Light Candidate filtering None
(Wang et al., 2025) IALP $\checkmark$ $\checkmark$ $\times$ $\times$ Medium (Few-shot+VLM) VLM None
(Agarwal and Sreepathy, 2024) TIC $\checkmark$ $\checkmark$ $\times$ $\times$ Light (Zero/Few-shot) None None
(de la Rosa et al., 2024) TRIP-PAL $\checkmark$ $\checkmark$ $\times$ $\times$ Light None None
(Guo et al., 2024) CaSTL $\checkmark$ $\checkmark$ $\times$ $\times$ Medium (Exp.) Python validator + LLM None
(Kalland, 2024) SemReBot2 $\checkmark$ $\checkmark$ $\times$ $\times$ Medium (Few-shot+Exp.) None None
(Kwon et al., 2024) $\checkmark$ $\checkmark$ $\times$ $\times$ Medium (Few-shot) None None
(Lee et al., 2024) PlanAID $\checkmark$ $\checkmark$ $\times$ $\times$ Light (Single-shot) Human deviation True
(Zhang et al., 2024a) PDDLEGO $\checkmark$ $\checkmark$ $\times$ $\times$ Medium (Few-shot+Exp.) LLM (PDDL-edit) None
(Zhang et al., 2024d) LaMMA-P $\checkmark$ $\checkmark$ $\times$ $\times$ Medium (Few-shot) FastDownward None
(Liu et al., 2023a) LLM+P $\checkmark$ $\checkmark$ $\times$ $\times$ Medium (Few-shot) None None
(Shirai et al., 2023) ViLaIn $\checkmark$ $\checkmark$ $\times$ $\times$ Medium (Few-shot+VLM) FastDownward None
(Birr et al., 2024) Auto-GPT+P $\times$ $\checkmark$ $\times$ $\times$ Light PDDL validator + Prolog True
(Grover and Mohan, 2024) $\times$ $\checkmark$ $\times$ $\times$ Medium (Few-shot) AI2Thor Env. None
(Izquierdo-Badiola et al., 2024) PlanCollabNL $\times$ $\checkmark$ $\times$ $\times$ Light LLM None
(Li et al., 2024b) SafePlanner $\times$ $\checkmark$ $\times$ $\times$ Medium (Exp.) None None
(Paulius et al., 2024) OLP $\times$ $\checkmark$ $\times$ $\times$ Light None None
(Singh et al., 2024b) DaTAPlan $\times$ $\checkmark$ $\times$ $\times$ Medium (Few-shot/CoT+Exp.) Human deviation True
(Singh et al., 2024a) TwoStep $\times$ $\checkmark$ $\times$ $\times$ Medium (Few-shot) LLM None
(Dagan et al., 2023) LLM+DP $\times$ $\checkmark$ $\times$ $\times$ Medium (Few-shot+Exp.) Alfworld Env. None
(Lyu et al., 2023) Faithful CoT $\times$ $\checkmark$ $\times$ $\times$ Light (CoT) None None
(Xie et al., 2023) $\times$ $\checkmark$ $\times$ $\times$ Medium (Few-shot) None None
(Collins et al., 2022) P+S $\times$ $\checkmark$ $\times$ $\times$ Medium (Few-shot) None None
Figure 4: Exp. = Explicit PDDL info. Feedback/Human Intervention provided at the level of the LLM-generated PDDL spec. itself. Papers reconstructed by L2P.
Appendix B Additional Information on PDDL
B.1 Why PDDL?
PDDL, as a declarative language, differs fundamentally from imperative languages like Python in how problems are expressed and solved. Declarative programming specifies what needs to be achieved rather than how to achieve it, leaving execution to an external planner. In contrast, imperative programming defines explicit step-by-step instructions, requiring precise control over execution flow. While LLMs struggle with logical reasoning in imperative programming due to its sequential dependencies, they perform well with declarative representations like PDDL. Using LLMs for PDDL model construction is not traditional code generation but rather knowledge structuring: organizing states, actions, and constraints into a formalized model. Instead of writing executable logic, LLMs assist in mapping natural language descriptions to structured symbolic representations.
B.2 PDDL Example
Demonstrated below is the Blocksworld PDDL domain file $\mathbb{DF}$:
(define (domain blocksworld)
(:requirements :strips)
(:predicates (clear ?x) (on-table ?x) (arm-empty) (holding ?x) (on ?x ?y))
(:action pickup
:parameters (?ob)
:precondition (and (clear ?ob) (on-table ?ob) (arm-empty))
:effect (and (holding ?ob) (not (clear ?ob)) (not (on-table ?ob)) (not (arm-empty)))
(define (problem blocksworld-problem)
(:domain blocksworld)
(:objects A B C) ; Blocks
(:init (ontable A) (ontable B) (on C A) (clear B) (clear C)) ; Initial state
(:goal (and (on A B) (on B C)))) ; Goal state
Unstack C from A
Put C on table
Pick up A
Stack A on B
Pick up B
Stack B on C
Appendix C (L2P) Framework
C.1 General Library Overview
L2P provides a comprehensive suite of tools for PDDL model creation and validation. The Builder classes enable users to prompt the LLM to generate essential components of a PDDL domain, such as types, predicates, and actions, along with their parameters, preconditions, and effects. It also supports task specification, including objects, initial, and goal states that correspond to the given domain. L2P features a customized Feedback Builder class that incorporates both LLM-generated feedback and human input, or a combination of the two. Additionally, the library includes a syntax validation tool that detects common PDDL syntax errors, using this feedback to improve the accuracy of the generated models—as illustrated in Figure 7, which shows how LLM feedback can be used to refine a PDDL problem specification.
C.2 Paper Reconstructions
L2P can recreate and encompass previous frameworks for converting natural language to PDDL, serving as a comprehensive foundation that integrates past approaches. L2P contains multiple paper reconstructions as examples of how existing methods can be implemented and compared within a unified system, demonstrating L2P’s flexibility and effectiveness in standardizing diverse NL-to-PDDL techniques. An example of Guan et al. (2023) “action-by-action” algorithm can be found in Figure 3; action and predicate output can be found in Figure 8.
C.3 L2P Usage Example
1 import os
2 from l2p.llm.openai import OPENAI
3 from l2p.utils import load_file
4 from l2p.domain_builder import DomainBuilder
5
6 domain_builder = DomainBuilder()
7
@@ C.3 L2P Usage Example
10
@@
14 predicates_prompt = load_file(f’{base_path}formalize_predicates.txt’)
@@
19 predicates, llm_output, validation_info = domain_builder.formalize_predicates(
@@
22 prompt_template=predicates_prompt,
23 types=types
@@
33 - (holding ?a - arm ?b - block) ; true if the arm ?a is currently holding the block ?b
34 - (on_table ?b - block) ; true if the block ?b is on the table
35 - (clear ?b - block) ; true if the block ?b is clear (no block on top of it)
36 - (on_top ?b1 - block ?b2 - block) ; true if the block ?b1 is on top of the block ?b2
@@
2 from l2p.utils.pddl_types import Predicate
3 from l2p.task_builder import TaskBuilder
4
@@
11 task_prompt = load_file(r’tests/usage/prompts/problem/formalize_task.txt’)
@@
17 objects, init, goal, llm_response, validation_info = task_builder.formalize_task(
@@
20 prompt_template=task_prompt,
@@
24
@@
27 domain_name="blocksworld",
28 problem_name="blocksworld_problem",
29 objects=objects,
30 initial=init,
31 goal=goal)
32
@@
41 arm1 - arm
42 blue_block - block
43 red_block - block
44 yellow_block - block
45 green_block - block
@@
53 (clear blue_block)
54 (clear green_block)
@@
58 (and
59 (on_top red_block green_block)
60 )
@@
2 from l2p.feedback_builder import FeedbackBuilder
@@
15 fb_pass, feedback_response = feedback_builder.task_feedback(
@@
22 types=types)
@@
26 [FEEDBACK]
27 ### JUDGMENT
28 ‘‘‘
29 My feedback on the provided PDDL problem file is as follows:
30